# Researcher Population Pyramids for Tracking Global Demographic and Gender Trajectories

### v1

@@ 4.1 Publication Data
Our analysis is based on OpenAlex, a large-scale, open bibliographic database containing hundreds of millions of publication records with extensive metadata across multiple disciplines [54]. We used its September 27, 2024 snapshot, from which we extracted 151,905,632 publications classified as ‘articles’ and published between 1950 and 2023. While our analysis focuses on this specific period (hereafter, $t_{\text{min}}=1950$ and $t_{\text{max}}=2023$), our framework is general and can be applied to any year range for which publication data are available. For each paper, we extracted its publication date and, for each author of the paper, their ID, full name, and affiliations. The author IDs are assigned by OpenAlex’s proprietary author-name disambiguation algorithm, which identified 64,665,774 unique authors in our dataset.
We assigned a primary country (or countries) of affiliation to each author $u$ as follows. For each author $u$, we first compiled a list of all country codes (ISO 3166-1 alpha-2) from their affiliations. We then computed the frequency distribution of these countries and assigned the country (or countries) with the highest frequency as the primary affiliation(s) for $u$. If multiple countries tied for the highest frequency, we assigned all of them to $u$. We excluded any author for whom no country was assigned. Through this process, we assigned at least one country to 34,507,112 authors. Throughout our manuscript, for a given country $c$, we refer to an author whose assigned countries include $c$ as an “author affiliated with country $c$.” If an author $u$ was assigned to multiple countries, they were included in the list of authors affiliated with each of their assigned countries.
@@ 4.2 Gender Assignment to Authors
For each author $u$ affiliated with country $c$, we extracted the first space-separated token from their full name (assumed to be the first name), and input it along with country $c$ into the CNB classifier to obtain an inferred gender $g$ and its associated probability $p_{g}$. We assigned gender $g$ to $u$ if $p_{g}\geq\theta_{c}$; otherwise, we left the gender unassigned. Using this procedure, we assigned binary gender labels to 15,104,205 authors affiliated with one of the 58 countries. Note that we excluded Lesotho, Papua New Guinea, and Somalia from our analysis of the 61 countries due to the small number of gender-assigned authors. Supplementary Table S7 reports the number of gender-assigned authors by country of affiliation. We acknowledge that this classification is based on a binary conception of gender (female or male) and does not capture the full spectrum of gender identities.
@@ 4.3 Inter-Publication Interval Threshold
We refine this approach by classifying each author as “active” or “inactive” in a given year, based on the survival function of inter-publication intervals. These intervals–the time between consecutive publications–are measured in years by dividing the number of days by 365 and are estimated separately for each country and gender. For each author who has published at least two papers, we extract their publication dates in chronological order to compute this sequence of intervals. For each country $c$ and gender $g$, we pool all such intervals and estimate the corresponding survival function. We then determine a threshold, $\Delta_{\text{IPI},c,g}$, defined as the largest inter-publication interval (in years) at which the survival probability remains above 1%. As detailed in Supplementary Section S4, these thresholds typically range from 7 to over 10 years.
@@ 4.4 Construction of Population Pyramids
For a given country $c$, binary gender $g$, and year $t\in[t_{\text{min}},t_{\text{max}}]$, we identify authors who satisfy a set of criteria adapted from the methodology of Huang et al. [18]. Specifically, we require that authors have: (i) published at least two papers by year $t$; (ii) the time elapsed from their first publication to their latest one as of year $t$, is between 1 and 40 years; and (iii) an annual productivity of less than 20 papers. To these established criteria, we introduce a novel fourth condition: (iv) they must be active in year $t$, based on the survival function of inter-publication intervals described in Section 4.3.
@@
To construct the population pyramid for a given country and year, we calculate the cumulative productivity as of that year for each active author affiliated with that country. This number, $k$, represents the length of the author’s most recent, uninterrupted sequence of publications. We determine $k$ by chronologically scanning an author’s entire publication record. We maintain a counter for the current continuous publication sequence. Starting with the first publication, this counter increases by one for each subsequent publication, as long as the time interval (in years) does not exceed the specified threshold ($\Delta_{\text{IPI},c,g}$). If an interval exceeds the threshold, the consecutive publication count ($k$) restarts at 1 with the next publication. The final value of the counter after scanning all publications represents the author’s cumulative productivity, $k$.
It is important to clarify that this procedure is not intended to suggest that a researcher’s prior work becomes irrelevant after a prolonged publication interval. Rather, it serves as a practical strategy to mitigate potential errors in author disambiguation within the OpenAlex database. Since author disambiguation remains a technically challenging task [64, 65, 53], publication records belonging to different individuals may occasionally be conflated. By treating sufficiently long intervals in publication as breaks in the publication record, we aim to minimize the impact of such errors on our population pyramid estimates.
@@
To project the population pyramid for country $c$ in years beyond $t_{\text{max}}$, we propagate the population pyramid at year $t_{\text{max}}$ using the transition observed between year $t_{\text{max}}-1$ and year $t_{\text{max}}$. We denote by $t^{\text{projection}}_{\text{max}}$ the end of the projection horizon beyond the observed period $[t_{\text{min}},t_{\text{max}}]$; we set $t^{\text{projection}}_{\text{max}}=2050$. Let $n_{k,c,g,t}$ denote the number of authors of gender $g$ affiliated with country $c$ whose cumulative productivity is $k$ at year $t$. For any $t_{\text{max}}+1\leq t\leq t^{\text{projection}}_{\text{max}}$, the projected population with cumulative productivity $k\geq 1$ evolves according to
@@
  • $a_{k,c,g,t}$ is the number of ‘newly active authors’ with a cumulative productivity of $k$ at year $t$ (i.e., those who have not published any papers between year $t-1-\Delta_{\mathrm{IPI},c,g}$ and year $t-1$ but whose cumulative productivity is $k$ at year $t$).
  • $P_{j\to k,c,g,t-1}$ is the probability that an author with a cumulative productivity of $j$ in year $t-1$ moves to a cumulative productivity of $k$ in year $t$.
@@
for all $t_{\text{max}}+1\leq t\leq t^{\text{projection}}_{\text{max}}$. We obtain every term on the right-hand side from the historical publication record. Specifically, we count $a_{k,c,g,t_{\text{max}}}$ for every $k\geq 1$, and we recursively compute $n_{k,c,g,t}$ for all $k\geq 1$ and $t_{\text{max}}+1\leq t\leq t^{\text{projection}}_{\text{max}}$. We estimate the transition probability $P_{j\to k,c,g,t_{\text{max}}-1}$ as the fraction of authors with a cumulative productivity of $j$ in $t_{\text{max}}-1$ who moved to a cumulative productivity of $k$ in $t_{\text{max}}$. We use these empirical quantities in Eq. (2) and iterate forward to obtain projected population pyramids for years beyond $t_{\text{max}}$. This stationarity assumption simplifies projection, though it may become less reliable over longer time horizons as changes in publishing-career patterns are not captured.
@@ Acknowledgments
This work was supported in part by ROIS-DS-JOINT (004RP2025). K.N. thanks the financial support by the JST ASPIRE Grant Number JPMJAP2328 and by JSPS KAKENHI Grant Numbers JP24K21056 and JP25H01122. T.M. thanks the financial support by JSPS KAKENHI Grant Numbers JP23H00042 and JP25K01458. T.M. is grateful to Daiki Takeda from the Faculty of Law, The University of Tokyo, for insightful discussions on the classifier.
@@
Publication and author data used in the analysis were extracted from the September 27, 2024, snapshot of OpenAlex [54]. The World Gender Name Dictionary 2.0 Dataset was obtained from the WIPO Knowledge Repository [2, 3]. The April 2024 snapshot of Moody’s Orbis database [1] was obtained for a fee.
@@
The code is available at https://github.com/kazuibasou/researcher_population_pyramids.
@@
Researcher Population Pyramids for Tracking Global Demographic and Gender Trajectories
@@
We established specific data requirements for constructing training, validation, and test sets from Orbis data. For each country included in our benchmarks, we aimed to create balanced validation and test sets, each comprising 450 female and 450 male executives. This specific sample size was selected primarily due to the limited data availability for Tunisia in the Orbis database, which recorded only 939 female executives. We ensured that the training set for each gender should be at least five times the size of the test set, necessitating a minimum of 2,250 individuals for training. Because our analysis has a particular focus on Arab countries, we accommodated countries with varying levels of data availability. This strategy involved a specific measure for data-scarce Arab nations: we supplemented their training data by incorporating names from other Arab countries that possessed more extensive datasets.
@@
To test the generalizability of our classifier beyond the Orbis data, we utilize the World Gender Name Dictionary (WGND) [2, 3]. The WGND ver. 2.0 contains 5,016,972 entries, each including a first name, binary gender, country, and the frequency of that name-gender combination within the country. These frequencies are primarily drawn from census and population registration data, though some sources reflect multi-year aggregations or alternative datasets [2].
@@
To prepare the data for our classifier, we first converted each first name into a feature vector. To ensure a consistent and comparable feature space across all countries, we tokenized every first name from the aggregated global training corpus into character-level $n$-grams. We set the maximum $n$-gram length to the length of the longest name within this global training corpus. We then transformed these tokenized names into high-dimensional vectors using a Term Frequency-Inverse Document Frequency (TF-IDF) weighting scheme, where each component represents the TF-IDF weight of a specific character $n$-gram.
@@
We evaluated our trained CNB classifier on both the Orbis and WGND test sets for each country by calculating the ROC AUC, F1 score, and the gender assignment rate (i.e., the fraction of gender-assigned test samples). For comparison, we also computed the same metrics for the CCT classifier. For all three metrics, higher values indicate better performance.
Table LABEL:table:s5 presents the benchmark results for both classifiers across all 61 countries in the Orbis test sets. As expected, the CNB classifier generally outperforms or matches the CCT classifier, given that the CNB model is directly trained on Orbis data. Across these 61 countries, the CNB classifier achieved (mean $\pm$ standard deviation) ROC AUCs of $0.979\pm 0.033$, F1 scores of $0.971\pm 0.032$, and gender assignment rates of $0.909\pm 0.098$. In comparison, the CCT classifier achieved ROC AUCs of $0.960\pm 0.048$, F1 scores of $0.970\pm 0.029$, and gender assignment rates of $0.847\pm 0.111$.
Table LABEL:table:s6 presents the benchmark results for both classifiers across all 61 countries in the WGND test sets. In contrast to Orbis, the CNB classifier performs comparably to or worse than the CCT classifier on the WGND test sets. Across these 61 countries, the CNB classifier achieved (mean $\pm$ standard deviation) ROC AUCs of $0.938\pm 0.049$, F1 scores of $0.942\pm 0.036$, and gender assignment rates of $0.747\pm 0.195$. In comparison, the CCT classifier achieved ROC AUCs of $0.987\pm 0.019$, F1 scores of $0.992\pm 0.012$, and gender assignment rates of $0.853\pm 0.065$.
This contrast highlights the complementary strengths of the two approaches: CNB effectively captures dataset-specific patterns, while CCT, built on broader cultural norms [6], generalizes well to population-level distributions. This observed difference also illustrates the inherent difficulty of inferring gender from names across countries and emphasizes the importance of cross-dataset evaluation. Unless we state otherwise, we use gender labels inferred and assigned via the CNB classifier for authors indexed in the OpenAlex.
@@
We obtained the estimated slopes $\{\beta_{c}\}_{c}$ for each of the nine Arab countries and the ten reference countries. Table S8 reports the estimated slope $\beta_{c}$ and the corresponding coefficient of determination $R^{2}$ for each country.
@@
We obtained the estimated slopes $\{\gamma_{c}\}_{c}$ for each of the nine Arab countries and the ten reference countries. Table S9 reports the estimated slope $\gamma_{c}$ and the corresponding coefficient of determination $R^{2}$ for each country.
S3.3 Change in Growth Rates Before and After the Arab Spring
This section details our methodology for examining whether the Arab Spring was associated with structural changes in researcher population growth. We specifically analyzed temporal trends in total author counts and in the proportion of female authors for Egypt and Tunisia. We compared estimated growth slopes before and after the Arab Spring using segmented linear regression.
To account for short-term volatility around the Arab Spring, we excluded the years 2010–2013 and split the time series into two segments: 2000–2009, which we termed the pre-Arab Spring period, and 2014–2023, the post-Arab Spring period.
We applied the same linear regression framework as described in the preceding subsections to estimate and compare these slopes across the two periods. For each period, we estimated the slope of the linear trend ($\hat{\beta}$ for author counts, $\hat{\gamma}$ for female author proportions). To account for potential autocorrelation in the time series data, we computed the standard errors of the slope estimates using Newey-West standard errors [7].
To test whether the slope differed significantly before and after the Arab Spring, we computed the following test statistic:
$t=\frac{\hat{\zeta}_{\mathrm{before}}-\hat{\zeta}_{\mathrm{after}}}{\sqrt{\mathrm{SE}_{\mathrm{before}}^{2}+\mathrm{SE}_{\mathrm{after}}^{2}}},$
where $\hat{\zeta}\in\{\hat{\beta},\hat{\gamma}\}$ and $\mathrm{SE}$ denotes the standard error of the slope estimate. Under the null hypothesis that the pre- and post-Arab Spring slopes are equal, the test statistic approximately follows a $t$-distribution with $10+10-4=16$ degrees of freedom, where 4 represents the total number of estimated parameters (two slopes and two intercepts) in the combined segmented regression model.
Tables S10 and S11 report the estimated slopes before and after the Arab Spring, along with the test statistics and $p$-values for the difference in slopes.
@@
Table LABEL:table:s12 shows the inter-publication interval threshold, $\Delta_{\mathrm{IPI},c,g}$, by country $c$ and gender $g$.
S5 Researcher Inflow and Gender Gap in Cumulative Productivity by Country in 2023
Table LABEL:table:s13 shows the researcher flow and the gender gap in cumulative productivity for each country in 2023.
S6 Compound Annual Growth Rate of the Number of Active Authors
@@
Austria $0.001$ $0.9600$
Bahrain $0.010$ $0.9910$
Belarus $0.001$ $0.9990$
@@
Russian Federation $0.001$ $0.9300$
Saudi Arabia $0.100$ $0.9300$
Serbia $0.001$ $0.9700$
Somalia $0.010$ $0.9000$
@@
Syrian Arab Republic $0.010$ $0.9930$
@@
Turkey $0.100$ $0.9000$
United Arab Emirates $0.001$ $0.9950$
@@
Table S5: Benchmark results on the Orbis test sets for the CNB and CCT classifiers. Columns AUC, F1, and GAT represent the ROC AUC, F1 score, and gender assignment rate, respectively.

### v2

@@ 4.1 Publication Data
Our analysis is based on OpenAlex, a large-scale, open bibliographic database containing hundreds of millions of publication records with extensive metadata across multiple disciplines [61]. We used its September 27, 2024 snapshot, from which we extracted 151,905,632 publications classified as “articles” and published between 1950 and 2023. While our analysis focuses on this specific period (hereafter, $t_{\text{min}}=1950$ and $t_{\text{max}}=2023$), our framework is general and can be applied to any year range for which publication data are available. For each paper, we extracted its publication date and, for each author of the paper, their ID, full name, and affiliations. The author IDs are assigned by OpenAlex’s proprietary author-name disambiguation algorithm.
We assigned a primary country (or countries) of affiliation to each author $u$ as follows. For each author $u$, we first compiled a list of all country codes (ISO 3166-1 alpha-2) from their affiliations. We then computed the frequency distribution of these countries and assigned the country (or countries) with the highest frequency as the primary affiliation(s) for $u$. If multiple countries tied for the highest frequency, we assigned all of them to $u$. We excluded any author for whom no country was assigned. Throughout our manuscript, for a given country $c$, we refer to an author whose assigned countries include $c$ as an “author affiliated with country cc.”
@@ 4.2 Gender Assignment to Authors
We thus began with the pool of 20,509,487 unique authors affiliated with one of the 61 countries. We first excluded 413,549 authors (2.0% of the pool) for whom a single primary country could not be uniquely determined (i.e., multiple countries tied for the highest frequency). For each remaining author $u$ affiliated with country $c$, we extracted the first space-separated token from their full name (assumed to be the first name), and input it along with country $c$ into the CNB classifier to obtain an inferred gender $g$ and its associated probability $p_{g}$. We assigned gender $g$ to $u$ if $p_{g}\geq\theta_{c}$; otherwise, we left the gender unassigned. Finally, we excluded three countries (Lesotho, Papua New Guinea, and Somalia) from our analysis due to the small number of gender-assigned authors. This process resulted in 14,745,796 gender-assigned authors affiliated with 58 countries. Supplementary Table S7 reports the number of gender-assigned authors by country of affiliation.
We acknowledge that our classification is based on a binary conception of gender (female or male) inferred from first names. This approach, while common in large-scale bibliometric analyses [15, 19], cannot account for individuals who identify as non-binary, transgender, or otherwise gender-diverse, and we recognize that it is an imperfect proxy for gender identity. Despite this limitation, we employ this approach as it remains one of the few feasible approaches for approximating gender demographics across countries.
@@ 4.3 Inter-Publication Interval Threshold
We refine this approach by classifying each author as “active” or “inactive” in a given year, based on the survival function of inter-publication intervals. These intervals—the time between consecutive publications—are measured in years by dividing the number of days by 365 and are estimated separately for each country and gender. For each author who has published at least two papers, we extract their publication dates in chronological order to compute this sequence of intervals. For each country $c$ and gender $g$, we pool all such intervals and estimate the corresponding survival function. We then determine a threshold, $\Delta_{\text{IPI},c,g}$, defined as the largest inter-publication interval (in years) at which the survival probability remains above 2%.
This 2% threshold was chosen after our preliminary analysis (see Supplementary Table S10), as we determined this value strikes a good balance between maximizing the inclusion of authors with diverse publication paces and ensuring the threshold duration remains plausible. A 1% threshold, for instance, yielded excessively long inter-publication intervals (e.g., 15 years or longer) for several countries such as Jamaica. Conversely, a 5% threshold resulted in intervals of five years or less for most countries, which could prematurely classify authors as inactive in disciplines with slower publication cycles.
@@ 4.4 Construction of Population Pyramids
For a given country $c$, binary gender $g$, and year $t\in[t_{\text{min}},t_{\text{max}}]$, we identify authors who satisfy a set of criteria extended from the methodology of Huang et al. [19]. Specifically, we require that authors meet four conditions: (i) they have published at least one paper by year $t$; (ii) the duration between their first publication and the end of year $t$ does not exceed 40 years; (iii) their total number of publications by year $t$ is at most 500, a threshold corresponding to approximately the top 0.1% of authors in our data; and (iv) they are classified as “active” in year $t$, based on the survival function of inter-publication intervals described in Section 4.3.
@@
To construct the population pyramid for a given country and year, we calculate the cumulative productivity as of that year for each active author affiliated with that country. This number, $k$, measures the continuity of an author’s recent research activity, defined as the length of the author’s most recent, uninterrupted sequence of publications. We determine $k$ by chronologically scanning an author’s entire publication record. We maintain a counter for the current continuous publication sequence. Starting with the first publication, this counter increases by one for each subsequent publication, as long as the time interval (in years) does not exceed the specified threshold ($\Delta_{\text{IPI},c,g}$). If an interval exceeds the threshold, the consecutive publication count ($k$) restarts at 1 with the next publication. It is important to clarify that this procedure is not intended to suggest that a researcher’s prior work becomes irrelevant after a prolonged publication interval. Rather, it serves as a practical strategy to mitigate potential errors in author disambiguation within the OpenAlex database. Since author disambiguation remains a technically challenging task [71, 72, 58], publication records belonging to different individuals may occasionally be conflated. By treating sufficiently long intervals in publication as breaks in the publication record, we aim to minimize the impact of such errors on our population pyramid estimates. The final value of the counter after scanning all publications represents the author’s cumulative productivity, $k$. We provide a supplementary analysis of cumulative productivity including its correlation with total productivity and publishing-career length in Supplementary Section S5.
@@
To project the population pyramid for country $c$ in years beyond $t_{\text{max}}$, we propagate the population pyramid at year $t_{\text{max}}$ using the transition observed between year $t_{\text{max}}-1$ and year $t_{\text{max}}$. We denote by $t^{\text{projection}}_{\text{max}}$ the end of the projection horizon beyond the observed period $t_{\text{min}},t_{\text{max}}]$; we set $t^{\text{projection}}_{\text{max}}=2050$. Let $n_{k,c,g,t}$ denote the number of authors of gender $g$ affiliated with country $c$ whose cumulative productivity is $k$ at year $t$. For any $t_{\text{max}}+1\leq t\leq t^{\text{projection}}_{\text{max}}$, the projected population with cumulative productivity $k\geq 1$ evolves according to
@@
  • $a_{k,c,g,t}$ is the number of ’newly active authors’ at year $t$. These are authors who were inactive at year t1t-1 but become active at year $t$ with a cumulative productivity of $k$.
  • $P_{j\to k,c,g,t-1}$ is the probability that an author with a cumulative productivity of $j$ in year t1t-1 moves to a cumulative productivity of $k$ in year $t$.
@@
for all $t_{\text{max}}+1\leq t\leq t^{\text{projection}}_{\text{max}}$. We obtain every term on the right‐hand side from the historical publication record. Specifically, we count $a_{k,c,g,t_{\text{max}}}$ for every $k\geq 1$, and we recursively compute $n_{k,c,g,t}$ for all $k\geq 1$ and $t_{\text{max}}+1\leq t\leq t^{\text{projection}}_{\text{max}}$. We estimate the transition probability $P_{j\to k,c,g,t_{\text{max}}-1}$ as the fraction of authors with a cumulative productivity of $j$ in $t_{\text{max}}-1$ who moved to a cumulative productivity of $k$ in $t_{\text{max}}$. We use these empirical quantities in Eq. (2) and iterate forward to obtain projected population pyramids for years beyond $t_{\text{max}}$. We emphasize that this analysis is based on a strong stationarity assumption regarding the inflow in $t_{\text{max}}$ and transition patterns observed from $t_{\text{max}}-1$ to $t_{\text{max}}$. This deterministic projection serves as a baseline scenario to illustrate potential long-term demographic trajectories if current conditions were to persist. This assumption is unlikely to hold over multi-decade horizons, as it does not account for cohort aging, policy shifts, or external shocks. The resulting projections should therefore be interpreted as stylized illustrations of current momentum.
@@ Acknowledgments
This manuscript was posted on a preprint: https://doi.org/10.48550/arXiv.2507.15500. We thank the anonymous reviewers for their constructive comments. This work was supported in part by ROIS-DS-JOINT (004RP2025). K.N. thanks the financial support by the JST ASPIRE Grant Number JPMJAP2328 and by JSPS KAKENHI Grant Numbers JP24K21056 and JP25H01122. T.M. thanks the financial support by JSPS KAKENHI Grant Numbers JP23H00042 and JP25K01458. T.M. is grateful to Daiki Takeda from the Faculty of Law, The University of Tokyo, for insightful discussions on the classifier.
@@
Publication and author data used in the analysis were extracted from the September 27, 2024, snapshot of OpenAlex [61]. The World Gender Name Dictionary 2.0 Dataset was obtained from the WIPO Knowledge Repository [2, 3]. The April 2024 snapshot of Moody’s Orbis database [1] was obtained for a fee. Access to Moody’s Orbis database was licensed and cannot be shared publicly. Information on how to request access is available at the following link: https://www.moodys.com/web/en/us/capabilities/company-reference-data/orbis.html.
@@
The code used for this manuscript is available at
https://github.com/kazuibasou/researcher_population_pyramids.
@@
Researcher Population Pyramids: Tracking Demographic and Gender Trajectories Across Countries
@@
We established specific data requirements for constructing training, validation, and test sets from Orbis data. For each country included in our benchmarks, we aimed to create balanced validation and test sets, each comprising 450 female and 450 male executives. This specific sample size was selected primarily due to the limited data availability for Tunisia in the Orbis database, which recorded only 939 female executives. We ensured that the training set for each gender was at least five times the size of the test set, necessitating a minimum of 2,250 individuals for training. Because our analysis has a particular focus on Arab countries, we accommodated countries with varying levels of data availability. This strategy involved a specific measure for data-scarce Arab nations: we supplemented their training data by incorporating names from other Arab countries that possessed more extensive datasets.
@@
To test the generalizability of our classifier beyond the Orbis data, we used the World Gender Name Dictionary (WGND) [2, 3]. The WGND ver. 2.0 contains 5,016,972 entries, each including a first name, binary gender, country, and the frequency of that name-gender combination within the country. These frequencies are primarily drawn from census and population registration data, though some sources reflect multi-year aggregations or alternative datasets [2].
@@
To prepare the data for our classifier, we first converted each first name into a feature vector. To ensure a consistent and comparable feature space across all countries, we tokenized every first name from the aggregated training corpus into character-level $n$-grams. We set the maximum $n$-gram length to the length of the longest name within this training corpus. We then transformed these tokenized names into high-dimensional vectors using a Term Frequency-Inverse Document Frequency (TF-IDF) weighting scheme, where each component represents the TF-IDF weight of a specific character $n$-gram.
@@
We evaluated our trained CNB classifier on both the Orbis and WGND test sets for each country by calculating the ROC AUC and F1 score. We also computed the gender assignment rate (GAR)—the fraction of gender-assigned test samples—separately for female and male samples ($\text{GAR}_{\text{f}}$ and $\text{GAR}_{\text{m}}$). For comparison, we computed the same set of metrics for the CCT classifier. For all these metrics, higher values indicate better performance.
Table LABEL:table:s5 presents the benchmark results for both classifiers across all 61 countries in the Orbis test sets. As expected, the CNB classifier generally outperforms or matches the CCT classifier, given that the CNB model is directly trained on Orbis data. Across these 61 countries, the CNB classifier achieved (mean $\pm$ standard deviation) ROC AUCs of $0.979\pm 0.034$, F1 scores of $0.973\pm 0.026$, and gender assignment rates of $0.896\pm 0.131$ for female and $0.907\pm 0.117$ for male test samples. In comparison, the CCT classifier achieved ROC AUCs of $0.960\pm 0.048$, F1 scores of $0.970\pm 0.029$, with corresponding gender assignment rates of $0.827\pm 0.137$ for female and $0.864\pm 0.112$ for male test samples.
Table LABEL:table:s6 presents the benchmark results for both classifiers across all 61 countries in the WGND test sets. In contrast to Orbis, the CNB classifier performs comparably to or worse than the CCT classifier on the WGND test sets. Across these 61 countries, the CNB classifier achieved (mean $\pm$ standard deviation) ROC AUCs of $0.938\pm 0.049$, F1 scores of $0.943\pm 0.036$, and gender assignment rates of $0.748\pm 0.198$ for female and $0.729\pm 0.211$ for male test samples. In comparison, the CCT classifier achieved ROC AUCs of $0.987\pm 0.020$, F1 scores of $0.992\pm 0.013$, with corresponding gender assignment rates of $0.845\pm 0.121$ for female and $0.839\pm 0.104$ for male test samples.
The two classifiers exhibit complementary strengths tied to the nature of the underlying datasets. The CCT classifier demonstrates superior generalization performance on the WGND benchmark, which is largely derived from general population data. In contrast, our CNB classifier, trained on the Orbis database of international executives, often outperforms the CCT classifier on the Orbis test set. This observed difference also illustrates the inherent difficulty of inferring gender from names across countries and emphasizes the importance of cross-dataset evaluation.
For the main analysis, we opted to use the CNB classifier. This decision was motivated by the specific nature of our target population: academic researchers. Researchers, much like the executives in the Orbis database, may constitute a highly mobile international population, a characteristic often termed “academic mobility” [7]. We reasoned that the demographic and naming patterns of this group might be more accurately captured by the Orbis data than by the more static, country-specific general population data represented by WGND. Therefore, we judged the CNB classifier to be a suitable choice for this specific analytical context.
@@
We obtained the estimated slopes $\{\beta_{c}\}_{c}$ for each of the nine Arab countries and the ten reference countries. To account for potential autocorrelation in the time-series data, we computed standard errors (SEs) and $p$-values using a Newey-West estimator. Table S8 reports the estimated slope $\beta_{c}$, its corresponding SE and $p$-value, and the coefficient of determination $R^{2}$ for each country.
@@
We obtained the estimated slopes $\{\gamma_{c}\}_{c}$ for each of the nine Arab countries and the ten reference countries. To account for potential autocorrelation in the time-series data, we computed standard errors (SEs) and $p$-values using a Newey-West estimator. Table S9 reports the estimated slope $\gamma_{c}$, its corresponding SE and $p$-value, and the coefficient of determination $R^{2}$ for each country.
@@
Table LABEL:table:s10 shows the inter-publication interval thresholds, $\Delta_{\mathrm{IPI},c,g}$, by country $c$ and gender $g$ when the survival probability threshold is set to 1%, 2%, and 5%.
S5 Analyses of Authors’ Career Metrics
We computed several career metrics for all authors active in 2023. In addition to our primary metric, cumulative productivity, we calculated the following for each author:
  • Total productivity: The total number of publications by the end of 2023.
  • Publishing-career length: The duration in years from their first publication to the end of 2023.
  • Publication gap experience: A binary indicator, true if the author has ever experienced an inter-publication interval longer than their corresponding $\Delta_{\text{IPI},c,g}$ threshold.
  • Returning author status: A binary indicator, true if the author was inactive in 2022 but became active in 2023, and had published at least one paper before 2023.
We first calculated the Spearman’s rank correlation between cumulative productivity, total productivity, and publishing-career length for each country and gender (see Table LABEL:table:s11). The Spearman’s rank correlation between cumulative productivity and total productivity was, as expected, very high, given that both are based on publication counts. However, the correlation was not perfect, which reflects that cumulative productivity is sensitive to the continuity of publication activity. In contrast, the correlation between cumulative productivity and publishing-career length was only moderate. This weaker correlation suggests that cumulative productivity is not a simple proxy for the length of publishing-career but rather captures an author’s recent publication momentum.
We further analyzed two metrics: the proportion of all active authors with a publication gap experience (see Table LABEL:table:s12), and the proportion of returning authors within newly active authors (see Table LABEL:table:s13). First, we observed a contrast between country groups. For both metrics, the proportions were substantially higher in the reference countries than in the Arab countries. On average, the proportion of authors with a publication gap experience was substantially higher in the reference countries (15.6% for female; 24.4% for male) than in the Arab countries (6.6% for female; 12.1% for male). Similarly, on average, the proportion of returning authors was substantially higher in reference countries (15.7% for female; 25.4% for male) than in Arab countries (5.7% for female; 9.5% for male). Second, we found a consistent gender difference across both metrics. In 56 out of 58 countries, a larger proportion of male authors had a publication gap experience compared to their female counterparts. Likewise, in 57 out of 58 countries, the proportion of returning authors was higher for men than for women. This counter-intuitive result is considered a consequence of our gender-specific inter-publication interval thresholds, which are empirically derived and are longer for women in most countries (48 out of 58 countries; see Table LABEL:table:s10).
S6 Researcher Inflow and Gender Gap in Cumulative Productivity by Country in 2023
Table LABEL:table:s14 shows the researcher flow and the gender gap in cumulative productivity for each country in 2023.
S7 Compound Annual Growth Rate of the Number of Active Authors
@@
S8 Analysis Across Research Domains
S8.1 Methods
We additionally conducted our analysis across the four high-level research domains defined by OpenAlex: “Health Sciences”, “Life Sciences”, “Physical Sciences”, and “Social Sciences”. Following the methodology used for country of affiliation (Section 4.1), we assigned research domains to each author based on the frequency of research domains associated with their publications. Specifically, we identified the research domain(s) that appeared most frequently across an author’s entire publication record. We retained only those authors whose research domain was uniquely determined (i.e., those with a single primary domain and no ties in frequency). For each combination of target year, country, and research domain, we constructed researcher population pyramids using the same criteria and methodology as those used to compute the researcher population pyramids shown in Fig. 2 in the main text.
S8.2 Results
Figures S1S4 present the researcher population pyramids for Egypt, Tunisia, Japan, and the United States across the four research domains for the years 2010, 2023, and 2050 (Health Sciences in Fig. S1, Life Sciences in Fig. S2, Physical Sciences in Fig. S3, and Social Sciences in Fig. S4). Our analysis reveals two key findings. First, the vertical reach of the pyramids varies across domains, reflecting distinct disciplinary publication norms; for example, researchers in the Health Sciences exhibit higher maximum levels of cumulative productivity, while those in the Social Sciences exhibit lower overall publication volumes. Second, despite these differences in vertical scale, the qualitative characteristics of the three research systems remain largely consistent within each country across all four domains. Egypt and Tunisia consistently exhibit the “Emerging” pattern; the United States exhibits the “Mature” pattern; and Japan exhibits the “Rigid” pattern in every domain analyzed. These results indicate that while disciplinary norms influence the absolute height (cumulative productivity levels) of the pyramids, our framework captures the demographic trajectories of national research ecosystems.
@@
Austria $0.001$ $0.9700$
Bahrain $0.010$ $0.9930$
Belarus $0.001$ $0.9992$
@@
Russian Federation $0.001$ $0.9400$
Saudi Arabia $0.100$ $0.9500$
Serbia $0.001$ $0.9600$
Somalia $0.010$ $0.9999$
@@
Syrian Arab Republic $0.100$ $0.9300$
@@
Türkiye $0.100$ $0.9000$
United Arab Emirates $0.001$ $0.9970$
@@
Table S5: Benchmark results on the Orbis test sets for the CNB and CCT classifiers. Columns AUC and F1 represent the ROC AUC and F1 score, respectively. $\text{GAR}_{\text{f}}$ and $\text{GAR}_{\text{m}}$ represent the gender assignment rates for the female and male test samples, respectively.